OpenStudy (anonymous):

John was graphing a function and noticed that at certain points, the graph reaches invisible lines the graph will never cross. Explain to John what the two types of invisible lines are and how to predict them. You may create your own example to aid in your reasoning. Use complete sentences. @kirbykirby

OpenStudy (anonymous):

How about this one

OpenStudy (anonymous):

When looking at a rational function, Jamal and Angie have two different thoughts. Jamal says that the function is defined at x = −3, x = −4, and x = 6. Angie says that the function is undefined at those x values. Describe a situation where Jamal is correct, and describe a situation where Angie is correct. Is it possible for a situation to exist where they are both correct? Justify your reasoning.

OpenStudy (kirbykirby):

The only thing that makes me think of "invisible lines" are possibly asymptotes for the 1st question

OpenStudy (anonymous):

No the second question

OpenStudy (anonymous):

Do you see it or nah

OpenStudy (kirbykirby):

When Angie is correct (undefined values at x=-3, -4, 6): \[f(x)=\frac{1}{(x+3)(x+4)(x-6)}\] When Jamal is correct (defined values for x=-3, -4, 6): \[f(x) = x \]

OpenStudy (kirbykirby):

You can't have a function when it is simultaneously defined and undefined for a certain value

OpenStudy (anonymous):

Oh

OpenStudy (kirbykirby):

does it make sense?

OpenStudy (anonymous):

Yes thank you